OpenStudy (anonymous):

lim s→6 [1/s - 1/6] / (s - 6)

OpenStudy (ksaimouli):

use the equation app

OpenStudy (ksaimouli):



OpenStudy (anonymous):

\[\lim_{s \rightarrow 6}\frac{ \left( \frac{ 1 }{ s } \right)-\left( \frac{ 1 }{ 6} \right) }{ s-6}\]

OpenStudy (anonymous):

Thank u i did @ksaimouli

OpenStudy (ksaimouli):